OpenStudy (anonymous):

Whenever you have a radical number, is it understood that it is positive or negative? For example is the square root of 7 understood to be the positive or negative square root of 7?

OpenStudy (turingtest):

when you take the square root of a number you have to leave open the possibility that the number itself is either positive or negative\[\sqrt{4}=\pm2\]because\[2*2=4\]and\[(-2)(-2)=4\]

OpenStudy (anonymous):

It's understood that it could be either one. There are two square roots of every number - a positive one and a negative one.

OpenStudy (anonymous):

okay, thank you

OpenStudy (mertsj):

It depends on how it is written. If you have x^2 = 4, there are two answers 2 and -2 but if you have  That symbol means the positive square root or 4. If you mean both of them you would have to put +/- in front of the radical.